OpenStudy (anonymous):

Integrate sin3xcos3x dx

OpenStudy (anonymous):

Let u = sin (3x) so du = 3 cos(3x) dx So your integral is of the form u du...the rest is straightforward. Add +C to your aanswer.

OpenStudy (anonymous):

caan u explain a bit furter

OpenStudy (anonymous):

If u = sin(3x), then the integral takes the form of 1/3 times the integral of u du Integral of u du is u^2/2...or (sin^2 (3x))/2 final answer is 1/3 times 1.2Sin^2(3x) = 1/6 sin^2(3x) + C